Solve the given word problem. You must use algebra and show your work.
There are two numbers where the smaller number is 2 less than the larger number. The total of three times the smaller number and twice the larger is forty-four. Find the numbers.

Answer by glitzgirl_14(11)

First you should set up 2 separate equations. I am going to use 's' for the smaller number and 'L' for the larger.
s = L - 2
3*s + 2*L = 44
Since you have the s equaling something, substitute 'L - 2' in for s in the second equation. So it would look like:
3( L - 2 ) + 2 * L = 44
Use the distributive property:
3L - 6 + 2L = 44
Put the L's together:
5L - 6 = 44
Bring the six over:
5L = 50
Divide by 5:
L= 10
Go back to the equation: s = L - 2
Substitute 10 in for L:
s = 10 - 2
s = 8
So the smaller number is 8 and the larger is 10.